### v1

# Recovery of rational functions via Hankel pencil method and sensitivities of the poles

Abstract
In the paper, we develop a new method for the recovery of rational functions. Our idea is based on the property that Fourier coefficients of rational functions have the exponential structure and reconstruction of this exponential structure with the ESPRIT method in the frequency domain. Further we present sensitivity analysis for poles of rational functions reconstructed with our method in case of unstructured and structured perturbations. Finally, we consider several numerical experiments and, using sensitivities, explain the recovery errors for poles.
Keywords: Rational functions, sparse exponential sums, matrix pencil, ESPRIT, Hankel matrices, sensitivity analysis, structured and unstructured perturbations.
AMS classification: 41A20, 42A16, 42C15, 65D15, 94A12.
1 Introduction
For $M\in\mathbb{N}$, we define a rational function $r(z)$ of type $(M-1,M)$ by means of its partial fraction decomposition
$r(z):=\sum\limits_{j=1}^{M}\frac{\gamma_{j}}{z-z_{j}},$ (1.1)
where $\gamma_{j},\,z_{j}\in\mathbb{C}$, $\gamma_{j}\neq 0$ and poles $z_{j}$ are pairwise distinct.
We study reconstruction of rational functions (1.1) using as input data a finite set of their Fourier coefficients. We use the property that Fourier coefficients of rational functions have the exponential structure and we reconstruct this exponential structure in the frequency domain applying the ESPRIT method [17, 15]. The idea of recovery of rational functions by estimation of signal parameters in the frequency domain is not new. Similar technique for the recovery of or approximation by rational functions was employed in [6, 18, 19, 20]. We present modification of these ideas, by applying the ESPRIT method for the exponential recovery. Since ESPRIT is based on using the Hankel pencil, it makes the method suitable for sensitivity analysis.
The other methods for the recovery of rational functions (in time domain) such as the AAA method [13], the RKFIT (Rational Krylov Fitting) method [3], the Vector Fitting algorithm [12] and the Loewner framework method [1], were considered in [11]. Note that, in [6, 18, 19, 20], similar to [11], reconstruction from given noisy data is limited to numerical observations.
In [21], for the Loewner framework method the authors study sensitivities of Loewner pencil eigenvalues (which are poles of a transfer function defined by the corresponding Loewner matrices) with respect to unstructured and structured perturbations. Following approach in [21], we present the sensitivity analysis for the poles computed by our method. According to our method, poles located inside and outside the unit circle are reconstructed separately as eigenvalues of some Hankel matrix pencils. Further we consider two kinds of Hankel pencil perturbations. For the sensitivity $\rho$ with respect to unstructured perturbations, we present the connection with condition numbers of Vandermonde matrices that are used in factorizations of the corresponding Hankel matrices. The sensitivity $\eta$ with respect to structured perturbations of Hankel pencil is studied in [22]. We present modification of these results for our special reconstruction problem. Note that the sensitivity of eigenvalues of Hankel pencil perturbations was studied also in [10], [2]. In [10] and [2], the authors measure both structured and unstructured perturbations by the same quantity. Following approach in [21], we consider two different quantities for structured and unstructured sensitivities respectively. Further using sensitivity analysis, we explain reconstruction errors for poles from noisy Fourier coefficients. We refer also to the paper [4] for further application of the sensitivity analysis of Hankel pencil.
The method offered in this paper is also very much related to our ESPIRA method for the exponential recovery [7][9]. In ESPIRA, we use the property that the classical Fourier coefficients of exponential sums have a rational structure and to reconstruct this rational structure we apply the AAA method for rational approximation [13] in the frequency domain. Thus the idea of the method for rational recovery used in this paper can be considered as converse to the idea of our ESPIRA method.
Outline. In Section 2, we present the main ideas of the ESPRIT method for exponential recovery. In Section 3 we consider sensitivity analysis for Hankel pencil eigenvalues. Subsection 3.1 is dedicated to unstructured perturbations and in Subsection 3.2, we consider structured perturbations. In Section 4, we describe our main results. In Subsection 4.1, we present main ideas of our method for the recovery of rational functions (1.1). Subsection 4.2 is dedicated to the sensitivity analysis of poles reconstructed with our method. Finally, in Section 5, we present numerical experiments.
Notation. Let us agree that throughout the paper, we will use the matrix notation $\mathbf{X}_{L,K}:=(X_{ij})_{i,j=1}^{L,K}$ for complex matrices of size $L\times K$ and the submatrix notation $\mathbf{X}_{L,K}(m:n,k:\ell)$, to denote a submatrix of $\mathbf{X}_{L,K}$ with rows indexed $m$ to $n$ and columns indexed $k$ to $\ell$, where first row and first column has index 1 (even though the row and column indices for the definition of the matrix may start with 0). For square matrices, we often use the short notation $\mathbf{X}_{K}$ instead of $\mathbf{X}_{K,K}$. By $\left(\mathbf{X}_{L,K}\right)^{\dagger}$ we denote the Moore-Penrose inverse of $\mathbf{X}_{L,K}$. We consider $\|\mathbf{X}_{L,K}\|_{2}:=\sigma_{\max}(\mathbf{X}_{L,K})$ the 2-norm or spectral norm, and $\|\mathbf{X}_{L,K}\|_{\mathrm{F}}:=\left(\sum\limits_{i=1}^{L}\sum\limits_{j=1}^{K}|X_{ij}|^{2}\right)^{1/2}=\left(\sum\limits_{i=1}^{\min\{L,K\}}\sigma_{i}^{2}\right)^{1/2}$ the Frobenius norm of the matrix $\mathbf{X}_{L,K}$. The $\ell_{1}$- and $\ell_{2}$-norm of the vector $\boldsymbol{x}=(x_{1},\ldots,x_{K})$ are given as usually by $\|\boldsymbol{x}\|_{1}=\sum\limits_{j=1}^{K}|x_{j}|$ and $\|\boldsymbol{x}\|_{2}=\left(\sum\limits_{j=1}^{K}|x_{j}|^{2}\right)^{1/2}$ respectively.
2 ESPRIT method for the recovery of exponential sums
As was described in Introduction, we will apply the ESPRIT method [17, 15] to reconstruct the exponential structure of the Fourier coefficients of rational functions (1.1). In this section, we explain the main ideas of this method.
Let $M\in\mathbb{N}$ and $f$ be an exponential sum defined by
$f(t):=\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{t},\ \gamma_{j},z_{j}\in\mathbb{C},\ \gamma_{j}\neq 0,$ (2.1)
where nodes $z_{j}$ are pairwise distinct. Let $L,N\in\mathbb{N}$ be given such that $M\leq L\leq N$, $L$ is the upper bound for $M$. We want to reconstruct an exponential sum (2.1), i.e. to reconstruct parameters $M$, $\gamma_{j}$ and $z_{j}$ for $j=1,\ldots,M$, using as input data samples $\boldsymbol{f}:=\big{(}f(0),f(1),\ldots,f(2N-1)\big{)}^{T}$.
For $n,m\in\mathbb{N}$ and a vector $\boldsymbol{a}:=(a_{1},\ldots,a_{m})^{T}$, we define a Vandermonde matrix $\mathbf{V}_{n,m}:=\mathbf{V}_{n,m}(\boldsymbol{a})=(a_{j}^{k})_{k=0,j=1}^{n-1,m}$. Let now $\boldsymbol{z}:=(z_{1},\ldots,z_{M})^{T}$ be a vector of nodes of the exponential sum (2.1) and let $\mathbf{\Gamma}_{M}:=\operatorname{diag}(\gamma_{j})_{j=1}^{M}$ and $\mathbf{Z}_{M}:=\operatorname{diag}(z_{j})_{j=1}^{M}$. We create a Hankel matrix ${\mathbf{H}}_{2N-L,L+1}\coloneqq\left(f(\ell+m)\right)_{\ell,m=0}^{2N-L-1,L}$ and consider its factorization
$\mathbf{H}_{2N-L,L+1}=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L+1,M}(\boldsymbol{z}).$ (2.2)
Since Vandermonde matrices in (2.2) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}$ has rank $M$. Therefore in case of exact given data, the order $M$ of the exponential sum (2.1) is reconstructed as the numerical rank of the matrix $\mathbf{H}_{2N-L,L+1}$.
Consider further a Hankel matrix pencil problem
$z\mathbf{H}_{2N-L,L}(0)-\mathbf{H}_{2N-L,L}(1),$ (2.3)
where $\mathbf{H}_{2N-L,L}(s):=\mathbf{H}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(\ell+k+s))_{k,\ell=0}^{2N-L-1,L-1}$ for $s=0,1$. Since the Hankel matrices $\mathbf{H}_{2N-L,L}(0)$ and $\mathbf{H}_{2N-L,L}(1)$ satisfy factorizations
$\mathbf{H}_{2N-L,L}(0)$ $=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}),$ (2.4)
$\mathbf{H}_{2N-L,L}(1)$ $=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}),$ (2.5)
we obtain that
$z\,{\mathbf{H}}_{2N-L,L}(0)-{\mathbf{H}}_{2N-L,L}(1)=\mathbf{V}_{2N-L,M}(\boldsymbol{z})\,\mathrm{diag}\left(z-z_{j}\right)_{j=1}^{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{L,M}(\boldsymbol{z}).$
From the last property, we conclude that nodes $z_{j}$, $j=1,\ldots,M$ in (2.1) are eigenvalues of the matrix pencil (2.3). Let us denote by $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ the vector of eigenvalues of (2.3). In case when $M$ is known, we can choose $N=L=M$ and compute $z_{j}$ as eigenvalues of
$z\mathbf{H}_{M}(0)-\mathbf{H}_{M}(1).$ (2.6)
Now we discuss how we solve the matrix pencil (2.3). Consider the SVD of ${\mathbf{H}}_{2N-L,L+1}$,
${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,L+1}{\mathbf{W}}_{L+1},$ (2.7)
where ${\mathbf{U}}_{2N-L}$ and ${\mathbf{W}}_{L+1}$ are unitary square matrices and, in the case of unperturbed data,
${\mathbf{D}}_{2N-L,L+1}=\left(\begin{array}[]{cc}\mathrm{diag}\left((\sigma_{j})_{j=1}^{M}\right)&{\mathbf{0}}\\ {\mathbf{0}}&{\mathbf{0}}\end{array}\right)\in{\mathbb{R}}^{(2N-L)\times(L+1)},$
where $\sigma_{1}\geq\ldots\geq\sigma_{M}$ denote the positive singular values of ${\mathbf{H}}_{2N-L,L+1}$. Thus (2.7) can be simplified to
${\mathbf{H}}_{2N-L,L+1}={\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1},$ (2.8)
where the (last) $L+1-M$ zero-columns in ${\mathbf{D}}_{2N-L,L+1}$ and the last $L+1-M$ rows in ${\mathbf{W}}_{L+1}$ are removed. Similarly as before, the SVDs of ${\mathbf{H}}_{2N-L,L}(0)$ and ${\mathbf{H}}_{2N-L,L}(1)$ can be directly derived from (2.8) and the matrix pencil in (2.3) now reads
$z\left({\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,1:L)$
$\hskip 142.26378pt-\left({\mathbf{U}}_{2N-L}\,{\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,2:L+1),$
or equivalently
$z\left({\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,1:L)-\left({\mathbf{D}}_{2N-L,M}{\mathbf{W}}_{M,L+1}\right)(1:2N-L,2:L+1).$
Multiplication from the left with $\left(\mathrm{diag}(\sigma_{j}^{-1})_{j=1}^{M},{\mathbf{0}}\right)\in{\mathbb{R}}^{M\times(2N-L)}$ yields the matrix pencil
$z\,{\mathbf{W}}_{M,L+1}(1:M,1:L)-{\mathbf{W}}_{M,L+1}(1:M,2:L+1).$ (2.9)
Thus, eigenvalues of the matrix pencil (2.3) can be computed as eigenvalues of the matrix pencil (2.9). Then we define ${\mathbf{W}}_{M,L}(0)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,1\!:\!L)$, ${\mathbf{W}}_{M,L}(1)\coloneqq{\mathbf{W}}_{L+1}(1\!:\!M,2\!:\!L+1)$, and determine the vector $\boldsymbol{z}=(z_{1},\ldots,z_{M})^{T}$ as vector of eigenvalues of $\left({\mathbf{W}}_{M,L}(0)^{T}\right)^{\dagger}{\mathbf{W}}_{M,L}(1)^{T}$.
Having nodes $z_{j}$, we reconstruct coefficients $\gamma_{j}$, $j=1,\ldots,M$ as a least squares solution of the system of linear equations
$\sum\limits_{j=1}^{M}\gamma_{j}z_{j}^{k}=f(k),\quad k=0,\ldots,2N-1,$
which can be also written in a matrix form
[…]
3 Sensitivities of Hankel pencil eigenvalues
In this section, we consider sensitivities of eigenvalues of the Hankel pencil (2.3). The concept of sensitivity arises from the perturbation of eigenvalues of the matrix pencil and for Hankel matrices was introduced in [10] and further considered in [2]. In [21], the notion of sensitivities was defined for Loewner matrix pencil. In the paper [21], the authors define two sensitivities, which correspond to structured and unstructured perturbations. Using similar ideas, in [22], the author considered the structured sensitivity of Hankel matrix pencil.
3.1 Sensitivity of Hankel pencil eigenvalues with unstructured perturbations
Following approach offered in [21] for Loewner pencil, we will show that in case of unstructured perturbations of Hankel pencil, the sensitivity of eigenvalues depends on the numerical conditions of Vandermonde matrices from factorizations (2.4) and (2.5) (for $N=L=M$). We consider the generalized eigenvalue problem (GEP) for the Hankel matrix pencil (2.6),
$\boldsymbol{p}^{T}\mathbf{H}_{M}(1)=z\boldsymbol{p}^{T}\mathbf{H}_{M}(0),\quad\mathbf{H}_{M}(1)\boldsymbol{q}=z\mathbf{H}_{M}(0)\boldsymbol{q},$ (3.1)
where $\boldsymbol{p}=(p_{1},\ldots,p_{M})^{T}$ and $\boldsymbol{q}=(q_{1},\ldots,q_{M})^{T}$ are right and left eigenvectors of the GEP (3.1) respectively. Then under the perturbation
$\tilde{\mathbf{H}}_{M}(0):=\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)},\ \ \tilde{\mathbf{H}}_{M}(1):=\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)},$
the first order approximation of the eigenvalue perturbation $z^{(1)}$ is given by
$z^{(1)}=\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}.$ (3.2)
Formula (3.2) is obtained from the perturbed system
$(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)$ $(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)})=(z+z^{(1)}+\ldots)(\boldsymbol{p}^{T}+(\boldsymbol{p}^{(1)})^{T}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{{\mathbf{H}_{M}(0)}}),$
$(\mathbf{H}_{M}(1)+\Delta_{\mathbf{H}_{M}(1)}))$ $(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots)=(z+z^{(1)}+\ldots)(\mathbf{H}_{M}(0)+\Delta_{\mathbf{H}_{M}(0)}))(\boldsymbol{q}+\boldsymbol{q}^{(1)}+\ldots),$
by retaining only first-order terms (see [10], [2], [21] and [22] for details).
Assume that the norm of perturbation matrices is bounded as follows
$\|\Delta_{\mathbf{H}_{M}(0)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(0)\|_{2},\ \ \|\Delta_{\mathbf{H}_{M}(1)}\|_{2}\leq\epsilon\|\mathbf{H}_{M}(1)\|_{2},$ (3.3)
where $\varepsilon>0$. Then from (3.2) taking into account (3.3), we get for the GEP (3.1)
$|z^{(1)}|$ $=\left|\frac{\boldsymbol{p}^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}_{L}}{\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}}\right|$
$\leq\frac{\|\boldsymbol{p}^{T}\|_{2}\,\|\Delta_{\mathbf{H}_{M}(1)}-z\,\Delta_{\mathbf{H}_{M}(0)}\|_{2}\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}\leq\epsilon\rho,$
where $\rho$ is the sensitivity of the eigenvalue $z$ of the GEP (3.1) defined by
$\rho:=\frac{\|\boldsymbol{p}^{T}\|_{2}\,(|z|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}\|_{2}}{|\boldsymbol{p}^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}|}.$ (3.4)
We derive further formulas for the exact computation and upper bounds of sensitivities $\rho_{j}$ of eigenvalues $z_{j}$, $j=1,\ldots,M$ of the GEP (3.1). Since Hankel matrices $\mathbf{H}_{M}(0)$ and $\mathbf{H}_{M}(1)$ satisfy factorizations (2.4) and (2.5) for $N=L=M$, left $\boldsymbol{p}^{(j)}$ and right $\boldsymbol{q}^{(j)}$ eigenvectors of the GEP (3.1) corresponding to the eigenvalue $z_{j}$ are given by formulas
$\boldsymbol{q}^{(j)}=\boldsymbol{p}^{(j)}=\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j},$ (3.5)
where $\boldsymbol{e}_{j}$ is the $j$th canonical vector in $\mathbb{R}^{M}$. Then using again the factorization (2.4) of the Hankel matrix $\mathbf{H}_{M}(0)$, we get
$(\boldsymbol{p}^{(j)})^{T}\,\mathbf{H}_{M}(0)\,\boldsymbol{q}^{(j)}$ $=\left(\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right)^{T}\,\mathbf{H}_{M}(0)\,\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}$
$=\boldsymbol{e}_{j}^{T}\,\mathbf{\Gamma}_{M}\,\boldsymbol{e}_{j}=\gamma_{j}.$ (3.6)
From (3.4) taking into account (3.6), we obtain that the sensitivity $\rho_{j}$ of the eigenvalue $z_{j}$ is given by
$\rho_{j}=\frac{1}{|\gamma_{j}|}\|(\boldsymbol{p}^{(j)})^{T}\|_{2}\,(|z_{j}|\,\|\mathbf{H}_{M}(0)\|_{2}+\|\mathbf{H}_{M}(1)\|_{2})\,\|\boldsymbol{q}^{(j)}\|_{2}.$ (3.7)
The ill-disposedness of the eigenvalue $z_{j}$ with respect to unstructured perturbations can be measured by checking whether $\rho_{j}$ is "large". Applying formulas (2.4), (2.5) and (3.5) to (3.7), we obtain for $j=1,\ldots,M$
$\rho_{j}$ $=\frac{1}{|\gamma_{j}|}\left\|\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\right\|_{2}^{2}\left(|z_{j}|\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}+\left\|\mathbf{V}_{M}(\boldsymbol{z})\,\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\,\mathbf{V}^{T}_{M}(\boldsymbol{z})\right\|_{2}\right).$ (3.8)
The formula (3.8) is used to compute the exact values of sensitivities $\rho_{j}$.
We proceed with estimation of (3.7). Since from (2.4) and (2.5), we have
$\|\mathbf{H}_{M}(0)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$
$\|\mathbf{H}_{M}(1)\|_{2}$ $\leq\|\mathbf{V}_{M}(\boldsymbol{z})\|_{2}\,\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}\,\|\mathbf{V}^{T}_{M}(\boldsymbol{z})\|_{2},$ (3.9)
and taking into account the fact $\|\mathbf{X}\,\boldsymbol{e}_{j}\|_{2}\leq\|\mathbf{X}\|_{2}$, we obtain
$\rho_{j}\leq\frac{|z_{j}|\,\|\mathbf{\Gamma}_{M}\|_{2}+\|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}}{|\gamma_{j}|}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$
where $\kappa(\mathbf{X}):=\|\mathbf{X}\|_{2}\,\|\mathbf{X}^{\dagger}\|_{2}$ denotes the condition number of the matrix $\mathbf{X}$. For diagonal matrices $\mathbf{\Gamma}_{M}$ and $\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}$, it holds
$\|\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|\gamma_{k}|,\ \ \ \|\mathbf{Z}_{M}\,\mathbf{\Gamma}_{M}\|_{2}=\max\limits_{k=1,\ldots,M}|z_{k}\gamma_{k}|.$
Thus
$\rho_{j}\leq\zeta_{j}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right),$ (3.10)
where $\zeta_{j}$ is defined by
$\zeta_{j}:=\frac{1}{|\gamma_{j}|}\left(|z_{j}|\max\limits_{k=1,\ldots,M}|\gamma_{k}|+\max\limits_{k=1,\ldots,M}|\gamma_{k}z_{k}|\right).$ (3.11)
Define further a vector of unstructured sensitivities $\boldsymbol{\rho}:=(\rho_{1},\ldots,\rho_{M})^{T}\in\mathbb{R}^{M}$ and let $\boldsymbol{\zeta}:=(\zeta_{1},\ldots,\zeta_{M})^{T}\in\mathbb{R}^{M}$. Then from (3.10) for the $\ell_{2}$-norm of the vector $\boldsymbol{\rho}$ we get
$\|\boldsymbol{\rho}\|_{2}\leq\|\boldsymbol{\zeta}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M}(\boldsymbol{z})\right).$
From (3.8), taking into account (3.9) and (3.11), we obtain
$\rho_{j}\leq\zeta_{j}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}\,\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}.$
By summing up the last inequalities over $j=1,\ldots,M$, we get
$\sum\limits_{j=1}^{M}\rho_{j}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\sum\limits_{j=1}^{M}\bigg{\|}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\,\boldsymbol{e}_{j}\bigg{\|}_{2}^{2}$
where $\zeta_{\max}:=\max_{j=1,\ldots,M}\zeta_{j}$. Finally using the property $\sum_{j=1}^{M}\|\mathbf{X}\boldsymbol{e}_{j}\|_{2}^{2}=\|\mathbf{X}\|_{\mathrm{F}}^{2}$, we obtain
$\|\boldsymbol{\rho}\|_{1}\leq\zeta_{\max}\big{\|}\mathbf{V}_{M}(\boldsymbol{z})\big{\|}_{2}^{2}\,\big{\|}\mathbf{V}^{\dagger}_{M}(\boldsymbol{z})\big{\|}_{\mathrm{F}}^{2}$
for the $\ell_{1}$-norm of the sensitivity vector $\boldsymbol{\rho}$.
3.2 Sensitivity Hankel pencil eigenvalues with structured perturbations
In this section we consider structured perturbations of the GEP (3.1) for the matrix pencil (2.6), i.e. the perturbation matrices have a particular structure, they are again Hankel matrices. We give a short description of the results obtained in [22]. Let $\boldsymbol{\delta}_{2M}:=(\delta_{0},\delta_{1},\ldots,\delta_{2M-1})^{T}$ be a vector of perturbations and we define
$\Delta_{\mathbf{H}_{M}(0)}:=(\delta_{k+\ell})_{k,\ell=0}^{M-1,M-1},\ \Delta_{\mathbf{H}_{M}(1)}:=(\delta_{k+\ell+1})_{k,\ell=0}^{M-1,M-1}.$
For the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$, $j=1,\ldots,M$ from (3.2) using (3.6), we obtain
$z^{(1)}_{j}=\frac{1}{\gamma_{j}}\,(\boldsymbol{p}^{(j)})^{T}\,(\Delta_{\mathbf{H}_{M}(1)}-z_{j}\,\Delta_{\mathbf{H}_{M}(0)})\,\boldsymbol{q}^{(j)}.$ (3.12)
Define further matrices $\mathbf{Q}^{(j)}_{M,2M}(0):=\left(Q^{(j)}_{k,m}(0)\right)_{k,m=1}^{M,2M}$ and $\mathbf{Q}^{(j)}_{M,2M}(1):=\left(Q^{(j)}_{k,m}(1)\right)_{k,m=1}^{M,2M}$ by their entries
$Q^{(j)}_{k,m}(0):=\begin{cases}q^{(j)}_{m-k+1},&1\leq m-k+1\leq M,\\ 0,&\text{otherwise},\end{cases}$
and
$Q^{(j)}_{k,m}(1):=\begin{cases}q^{(j)}_{m-k},&1\leq m-k\leq M,\\ 0,&\text{otherwise}.\end{cases}$
Then
$\Delta_{\mathbf{H}_{M}(0)}\,\boldsymbol{q}^{(j)}=\mathbf{Q}_{M,2M}^{(j)}(0)\,\boldsymbol{\delta}_{2M},\ \Delta_{\mathbf{H}_{M}(1)}\,\boldsymbol{q}^{(j)}=\mathbf{Q}_{M,2M}^{(j)}(1)\,\boldsymbol{\delta}_{2M}.$
Since for $m=1,\ldots,2M$
$\left((\boldsymbol{p}^{(j)})^{T}\,\mathbf{Q}^{(j)}_{M,2M}(0)\right)_{m}$ $=\sum\limits_{k=1}^{2M}p^{(j)}_{k}\,Q^{(j)}_{k,m}(0)=\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}p^{(j)}_{k}\,q^{(j)}_{m-k+1},$
$\left((\boldsymbol{p}^{(j)})^{T}\,\mathbf{Q}_{M,2M}^{(j)}(1)\right)_{m}$ $=\sum\limits_{k=1}^{M}p^{(j)}_{k}\,Q_{k,m}^{(j)}(1)=\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k},$
taking into account (3.12), we define vectors $\boldsymbol{\mathcal{S}}^{(j)}_{2M}:=\left(S^{(j)}_{1},\ldots,S^{(j)}_{2M}\right)^{T}\in\mathbb{C}^{2M}$, $j=1,\ldots,M$ by their components
$S^{(j)}_{m}:=\frac{1}{\gamma_{j}}\left(\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}p^{(j)}_{k}\,q^{(j)}_{m-k}-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}p^{(j)}_{k}\,q^{(j)}_{m-k+1}\right),$
$m=1,\ldots,2M$. Then the first order approximation of the eigenvalue perturbation $z_{j}^{(1)}$ in (3.12) can be given by
$z_{j}^{(1)}=(\boldsymbol{\mathcal{S}}^{(j)}_{2M})^{T}\,\boldsymbol{\delta}_{2M},\quad j=1,\ldots,M.$ (3.13)
The component $S^{(j)}_{m}$ shows the influence of the perturbation of the $m$th measurement (in this case it is the function value $f(m-1)$) on the $j$th eigenvalue $z_{j}$ of the Hankel matrix pencil (2.6). Using (3.5), we rewrite components $S^{(j)}_{m}$ in the following way
$S^{(j)}_{m}=$ $\frac{1}{\gamma_{j}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M\}}^{\min\{M,m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k,j}$
$-z_{j}\,\sum\limits_{k=\max\{1,m-M+1\}}^{\min\{M,m\}}\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{k,j}\,\bigg{(}\left(\mathbf{V}^{T}_{M}(\boldsymbol{z})\right)^{\dagger}\bigg{)}_{m-k+1,j}\bigg{)}.$
Formula (3.13) can be also written in a matrix form
$\boldsymbol{z}^{(1)}=\mathbf{\mathcal{S}}_{M,2M}\,\boldsymbol{\delta}_{2M},$
where $\boldsymbol{z}^{(1)}:=(z_{1}^{(1)},\ldots,z_{M}^{(1)})^{T}$ and $\mathbf{\mathcal{S}}_{M,2M}:=(S_{j,m})_{j,m=1}^{M,2M}\in\mathbb{C}^{M\times 2M}$ be a matrix with the entries $S_{j,m}:=S^{(j)}_{m}$ that quantify the perturbation of the eigenvalues of matrix pencil (2.6).
Finally, we define a structured sensitivity of the eigenvalue $z_{j}$ of matrix pencil (2.6) with respect to the $m$th measurement by
$\eta_{j,m}:=|S_{j,m}|$
for $j=1,\ldots,M$ and $m=1,\ldots,2M$. Then the eigenvalue sensitivity $\eta_{j}$ for the eigenvalue $z_{j}$ with respect to structured perturbations of Hankel matrix pencil (2.6) is defined as the $\ell_{2}$-norm of the vector $\boldsymbol{\mathcal{S}}^{(j)}_{2M}\in\mathbb{C}^{2M}$. Thus we further introduce a vector $\boldsymbol{\eta}:=(\eta_{1},\ldots,\eta_{M})^{T}\in\mathbb{R}^{M}$ with its entries
$\eta_{j}:=\|\boldsymbol{\mathcal{S}}^{(j)}_{2M}\|_{2}=\sqrt{\eta_{j,1}^{2}+\ldots+\eta^{2}_{j,2M}},\quad j=1,\ldots,M.$
The ill-disposedness of the eigenvalue $z_{j}$ with respect to structured perturbations can be measured by checking whether $\eta_{j}$ is "large". The entry $\eta_{j}$ determines also the standard deviation of the eigenvalue perturbation $z_{j}^{(1)}$ (see [22] for details).
4 Recovery of rational functions via Hankel matrix pencil method and sensitivities of the poles
In this section, we describe the main ideas of our new method for the rational reconstruction and present analysis of the sensitivities of the poles.
4.1 Description of the method
Let $r(t)$ be a rational function defined by (1.1) and by $\mathbb{D}$ we denote the unit circle. Following approach in [19], for $k\in\mathbb{Z}$ we consider the integral
$\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t},$
which is closely related to the classical Fourier coefficient of a function $r(\mathrm{e}^{\mathrm{i}\varphi})$,
$\widehat{r}_{k}:=\frac{1}{2\pi}\int\limits_{0}^{2\pi}r(\mathrm{e}^{\mathrm{i}\varphi})\mathrm{e}^{-\mathrm{i}k\varphi}\,\mathrm{d}\varphi=\frac{1}{2\pi\mathrm{i}}\int\limits_{\mathbb{D}}\frac{r(t)}{t^{k}}\,\frac{\mathrm{d}t}{t}.$
Then (see, for example, [19])
$\widehat{r}_{k}=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{-(k+1)},&k\leq-1,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{-(k+1)},&k\geq 1.\end{cases}$ (4.1)
From (4.1), we conclude that Fourier coefficients $\widehat{r}_{k}$ with negative indeces $k=-1,-2,\ldots$ contain information about poles inside the unite circle, and Fourier coefficients $\widehat{r}_{k}$ with positive indises $k=1,2,\ldots$ contain information about poles outside the unite circle.
Taking into account (4.1), we define an exponential sum $f(t)$ by
$f(t)=\begin{cases}\sum\limits_{|z_{j}|<1}\gamma_{j}z_{j}^{t},&t\geq 0,\\ -\sum\limits_{|z_{j}|>1}\gamma_{j}z_{j}^{t},&t\leq-2,\end{cases}$ (4.2)
and from (4.1) we obtain the following interpolation property
$f(-(k+1))=\widehat{r}_{k},\quad k\in\mathbb{Z}\setminus\{0\}.$ (4.3)
Therefore the rational reconstruction problem (1.1) can be reformulated as the exponential interpolation problem (4.3). Our further goal is to reconstruct an exponential sum (4.2), using samples $f(m)$, $m\in\mathbb{Z}\setminus\{-1\}$ as input data. From (4.2), we get that values $f(m)$ with $m=0,1,2,\ldots$ contain information about nodes $z_{j}$ inside $\mathbb{D}$, and values $f(m)$ with $m=-2,-3,\ldots$ contain information about nodes $z_{j}$ outside $\mathbb{D}$.
Let $M_{1}$ and $M_{2}$ be numbers of nodes $z_{j}$ of the exponential sum (4.2) located inside $\mathbb{D}$ and outside $\mathbb{D}$ respectively. We assume that a number of poles $M$ as well as $M_{1}$ and $M_{2}$ are also unknown. Let $L,N\in\mathbb{N}$ and $M_{1},M_{2}\leq L\leq N$, where $L$ is the upper bound for $M$. We use values $f(k)$, $k=0,\ldots,2N-1$ ($\widehat{r}_{-k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ inside $\mathbb{D}$, and values $f(-k)$, $k=2,\ldots,2N+1$ ($\widehat{r}_{k}$, $k=1,\ldots,2N$) to reconstruct nodes $z_{j}$ outside $\mathbb{D}$. Further we rewrite an exponential sum (4.2) in the following way
$f(t)=\begin{cases}\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{t},&t\geq 0,\\ -\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{t}&t\leq-2,\end{cases}$ (4.4)
i.e. by $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ we denote poles located inside $\mathbb{D}$ and let $\gamma^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ be the corresponding coefficients and by $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ we denote poles located outside $\mathbb{D}$ and let $\gamma^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$ be the corresponding coefficients.
To reconstruct exponential sums (4.4), we use the ESPRIT method described in Section 2. We present some modification of this method for our special parameter estimation problem (4.4). Note that the ESPRIT is not the only method to apply in the frequency domain. We can use the other methods for exponential recovery, for example, MPM (matrix-pencil method) [15] or approximate Prony method [16]. In [19], the author uses the AAK theory [14] in the frequency domain to reconstruct (1.1) and a matrix-valued version of (1.1).
Let $\boldsymbol{z}^{(-)}\in\mathbb{C}^{M_{1}}$ and $\boldsymbol{z}^{(+)}\in\mathbb{C}^{M_{2}}$ be two vectors defined by
$\boldsymbol{z}^{(-)}:=(z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}),\ \ \boldsymbol{z}^{(+)}:=\left(\frac{1}{z_{j_{1}}^{(+)}},\ldots,\frac{1}{z^{(+)}_{j_{M_{2}}}}\right).$
Let us define also diagonal matrices
$\mathbf{\Gamma}^{(-)}_{M_{1}}:=\mathrm{diag}(\gamma^{(-)}_{j_{n}})_{n=1}^{M_{1}},\ \mathbf{\Gamma}^{(+)}_{M_{2}}:=\mathrm{diag}(\gamma^{(+)}_{j_{n}})_{n=1}^{M_{2}},\ \mathbf{Z}_{M_{1}}^{(+)}:=\mathrm{diag}\left(z^{(-)}_{j_{n}}\right)_{n=1}^{M_{1}},\ \mathbf{Z}_{M_{2}}^{(+)}:=\mathrm{diag}\left(\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}.$
According to the ESPRIT method described in Section 2, to reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the unit circle, we use the Hankel matrix
$\mathbf{H}^{(-)}_{2N-L,L+1}:=\left(f(\ell+k)\right)_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{-(k+\ell)}\right)_{k=0,\ell=1}^{2N-L-1,L+1},$
and to reconstruct nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the unit circle, we use the Hankel matrix
$\mathbf{H}^{(+)}_{2N-L,L+1}:=(f(-(\ell+k+2)))_{k,\ell=0}^{2N-L-1,L}=\left(\widehat{r}_{k+\ell}\right)_{k=0,\ell=1}^{2N-L-1,L+1}.$
Consider factorizations of these matrices
$\mathbf{H}^{(-)}_{2N-L,L+1}$ $=\mathbf{V}_{2N-L,M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}^{(-)}_{M_{1}}\,\mathbf{V}^{T}_{L+1,M_{1}}(\boldsymbol{z}^{(-)}),$ (4.5)
$\mathbf{H}^{(+)}_{2N-L,L+1}$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L+1,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.6)
Since all Vandermonde matrices in (4.5) and (4.6) have full column ranks, the matrix $\mathbf{H}_{2N-L,L+1}^{(-)}$ has rank $M_{1}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ has rank $M_{2}$. Therefore in case of exact given data, the numbers $M_{1}$ and $M_{2}$ are reconstructed as numerical ranks of matrices $\mathbf{H}_{2N-L,L+1}^{(-)}$ and $\mathbf{H}_{2N-L,L+1}^{(+)}$ respectively.
First we consider the reconstruction of nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$, located inside the $\mathbb{D}$. Since in this case, the Hankel matrices $\mathbf{H}^{(-)}_{2N-L,L+1}$ have structure similar to the structure of the Hankel matrices $\mathbf{H}_{2N-L,L+1}$ in the Section 2, following ideas of the Section 2, we get that nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ are reconstructed as eigenvalues of the Hankel matrix pencil
$z\mathbf{H}^{(-)}_{2N-L,L}(0)-\mathbf{H}^{(-)}_{2N-L,L}(1),$ (4.7)
where for $s=0,1$
$\mathbf{H}^{(-)}_{2N-L,L}(s):=\mathbf{H}^{(-)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(\ell+k+s))_{k,\ell=0}^{2N-L-1,L-1}.$
Consider now the case of nodes $z^{(+)}_{j_{n}}$, $n=1,\ldots,M_{2}$, located outside the $\mathbb{D}$. For $s=0,1$, define matrices
$\mathbf{H}^{(+)}_{2N-L,L}(s):=\mathbf{H}^{(+)}_{2N-L,L+1}(1:2N-L,1+s:L+s)=(f(-(\ell+k+s+2)))_{k,\ell=0}^{2N-L-1,L-1},$
and consider their factorizations
$\mathbf{H}^{(+)}_{2N-L,L}(0)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$ (4.8)
$\mathbf{H}^{(+)}_{2N-L,L}(1)$ $=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}).$ (4.9)
Since according to (4.8) and (4.9) we have
$z\,{\mathbf{H}}^{(+)}_{2N-L,L}(0)-{\mathbf{H}}^{(+)}_{2N-L,L}(1)=-\mathbf{V}_{2N-L,M_{2}}(\boldsymbol{z}^{(+)})\,\mathrm{diag}\left(z-\frac{1}{z^{(+)}_{j_{n}}}\right)_{n=1}^{M_{2}}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{L,M_{2}}(\boldsymbol{z}^{(+)}),$
then inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil
$z\mathbf{H}^{(+)}_{2N-L,L}(0)-\mathbf{H}^{(+)}_{2N-L,L}(1).$ (4.10)
If we assume that $M_{1}$ and $M_{2}$ are known, we can reconstruct nodes $z^{(-)}_{j_{n}}$, $n=1,\ldots,M_{1}$ as eigenvalues of the Hankel matrix pencil ($N=L=M_{1}$ in (4.7))
$z\mathbf{H}^{(-)}_{M_{1}}(0)-\mathbf{H}^{(-)}_{M_{1}}(1),$ (4.11)
and inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of nodes $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ are reconstructed as eigenvalues of the Hankel matrix pencil ($N=L=M_{2}$ in (4.10))
$z\mathbf{H}^{(+)}_{M_{2}}(0)-\mathbf{H}^{(+)}_{M_{2}}(1).$ (4.12)
Having nodes $z^{(-)}_{j_{n}}$ and $z^{(+)}_{j_{n}}$, coefficients $\gamma^{(-)}_{j_{n}}$ and $\gamma^{(+)}_{j_{n}}$ can be computed as least squares solutions of the overdetermined systems of linear equations
$\sum\limits_{n=1}^{M_{1}}\gamma^{(-)}_{j_{n}}(z^{(-)}_{j_{n}})^{-(k+1)}=\widehat{r}_{-k},\quad k=1,\ldots,2N,$
and
$-\sum\limits_{n=1}^{M_{2}}\gamma^{(+)}_{j_{n}}(z^{(+)}_{j_{n}})^{-(k+1)}=\widehat{r}_{k},\quad k=1,\ldots,2N,$
respectively.
Remark 4.1.
Note that coefficients $\widehat{r}_{k}$ can be also computed applying the quadrature formulas on some grid of points. In that case, we will use the function values instead of Fourier coefficients as input data for the algorithm. This question was discussed in details in [19], where the author consider the random access model and the Matsubara model as a set of grid points for the evaluation of the transform $\widehat{r}_{k}$. In this paper, we are interested in the presentation of a modification of the method from [19] and in the explanation of the results of this method regarding recovery of noisy data with the use of sensitivities.
4.2 Sensitivities of poles
In this subsection, we consider sensitivities of poles $z_{j}$ of the reconstruction problem (1.1) that is solved with the method described in Subsection 4.1. First we consider the case of unstructured perturbations. Poles $z_{j_{n}}^{(-)}$, $n=1,\ldots,M_{1}$ located inside the unit circle $\mathbb{D}$, are reconstructed as eigenvalues of the matrix pencil (4.11). Since Hankel matrices in the matrix pencil (4.11) have similar structure as Hankel matrices in the matrix pencil (2.6), we apply ideas from Subsection 3.1 without any changes and obtain for the sensitivities $\rho_{n}^{(-)}$, $n=1,\ldots,M_{1}$ of the poles $z^{(-)}_{j_{1}},\ldots,z^{(-)}_{j_{M_{1}}}$ in the $\mathbb{D}$ the formula for the exact computation
$\rho_{n}^{(-)}$ $=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left\|\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right\|_{2}^{2}$
×$\times\left(|z_{j_{n}}^{(-)}|\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}+\left\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\,\mathbf{Z}_{M_{1}}\,\mathbf{\Gamma}_{M_{1}}\,\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right\|_{2}\right),$ (4.13)
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{1}}$ and the upper bound
$\rho_{n}^{(-)}\leq\zeta_{n}^{(-)}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),$ (4.14)
with
$\zeta_{n}^{(-)}:=\frac{1}{\left|\gamma_{j_{n}}^{(-)}\right|}\left(\left|z_{j_{n}}^{(-)}\right|\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}\right|+\max\limits_{k=1,\ldots,M_{1}}\left|\gamma_{j_{k}}^{(-)}z_{j_{k}}^{(-)}\right|\right).$
Now consider the case of the sensitivities $\rho_{n}^{(+)}$ of poles $z^{(+)}_{j_{n}}$ for $n=1,\ldots,M_{2}$ located outside the unit circle $\mathbb{D}$. Inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of poles $z^{(+)}_{j_{n}}$ are reconstructed as eigenvalues of the matrix pencil (4.12). That means that we will compute sensitivities $\rho_{n}^{(+)}$ of the inverse $\frac{1}{z^{(+)}_{j_{n}}}$ of the poles $z^{(+)}_{j_{n}}$, but we will say that $\rho_{n}^{(+)}$ are also sensitivities of poles $z^{(+)}_{j_{n}}$.
To present formulas for sensitivities $\rho_{n}^{(+)}$, we have to apply some modifications to the results of the Subsection 3.1. Matrices $\mathbf{H}^{(+)}_{M_{2}}(0)$ and $\mathbf{H}^{(+)}_{M_{2}}(1)$ satisfy factorizations
$\mathbf{H}^{(+)}_{M_{2}}(0)=-\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)}),$ (4.15)
$\mathbf{H}^{(+)}_{M_{2}}(1)=-\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)}).$ (4.16)
Thus left and right eigenvectors of the GEP of the Hankel matrix pencil (4.12) corresponding to the eigenvalue $\frac{1}{z^{(+)}_{j_{n}}}$ are given by formulas
$\boldsymbol{q}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},\quad\boldsymbol{p}^{(n)}=\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n},$ (4.17)
where $\boldsymbol{e}_{n}$ is the $n$th canonical (column) vector in $\mathbb{R}^{M_{2}}$ Taking into account (4.15) and (4.17), we get
$(\boldsymbol{p}^{(n)})^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\boldsymbol{q}^{(n)}$ $=\left(\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\right)^{T}\,\mathbf{H}^{(+)}_{M_{2}}(0)\,\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}$
$=-\boldsymbol{e}_{n}^{T}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,\left(\mathbf{Z}_{M_{2}}^{(+)}\right)^{2}\,\boldsymbol{e}_{n}=-\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}.$ (4.18)
From (3.4) applying (4.15)–(4.18), we obtain the formula for the exact computation of sensitivities $\rho_{n}^{(+)}$ of the poles $z^{(+)}_{j_{n}}$ (the eigenvalues $\frac{1}{z^{(+)}_{j_{n}}}$ of the matrix pencil (4.12)) located outside the unit circle
$\rho_{n}^{(+)}=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z_{j_{n}}^{(+)})^{-2}|}\bigg{\|}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\,\boldsymbol{e}_{n}\bigg{\|}_{2}^{2}$
×$\times\bigg{(}\frac{1}{|z_{j_{n}}^{(+)}|}\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{2}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}+\left\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\,\mathbf{\Gamma}^{(+)}_{M_{2}}\,(\mathbf{Z}_{M_{2}}^{(+)})^{3}\,\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right\|_{2}\bigg{)},$ (4.19)
and the upper bound
$\rho^{(+)}_{n}\leq\frac{\frac{1}{|z_{j_{n}}^{(+)}|}\,\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}+\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$
Now since
$\big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{2}\mathbf{\Gamma}_{M_{2}}^{(+)}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|,\ \ \ \big{\|}\big{(}\mathbf{Z}^{(+)}_{M_{2}}\big{)}^{3}\,\mathbf{\Gamma}^{(+)}_{M_{2}}\big{\|}_{2}=\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|,$
we get for $n=1,\ldots,M_{2}$
$\rho_{n}^{(+)}\leq\zeta_{n}^{(+)}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right),$ (4.20)
with
$\zeta_{n}^{(+)}:=\frac{1}{|\gamma_{j_{n}}^{(+)}\,(z^{(+)}_{j_{n}})^{-2}|}\left(\frac{1}{|z_{j_{n}}^{(+)}|}\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-2}|+\max\limits_{k=1,\ldots,M_{2}}|\gamma^{(+)}_{j_{k}}\,(z^{(+)}_{j_{k}})^{-3}|\right).$
Define further two vectors of unstructured sensitivities $\boldsymbol{\rho}^{(-)}:=(\rho_{1}^{(-)},\ldots,\rho^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\rho}^{(+)}:=(\rho_{1}^{(+)},\ldots,\rho^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$, and let $\boldsymbol{\zeta}^{(-)}:=(\zeta_{1}^{(-)},\ldots,\zeta^{(-)}_{M_{1}})^{T}\in\mathbb{R}^{M_{1}}$ and $\boldsymbol{\zeta}^{(+)}:=(\zeta_{1}^{(+)},\ldots,\zeta^{(+)}_{M_{2}})^{T}\in\mathbb{R}^{M_{2}}$. Then using (4.14) and (4.20), we write the following upper bounds for the $\ell_{2}$-norm of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{2}\leq\|\boldsymbol{\zeta}^{(-)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\right),\quad\|\boldsymbol{\rho}^{(+)}\|_{2}\leq\|\boldsymbol{\zeta}^{(+)}\|_{2}\,\kappa^{2}\left(\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\right).$ (4.21)
Using ideas similar as in Subsection 3.1, we write also the upper bounds for the $\ell_{1}$-norm of the sensitivity vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ $\leq\zeta^{(-)}_{\max}\,\|\mathbf{V}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{1}}(\boldsymbol{z}^{(-)})\|_{\mathrm{F}}^{2}$ (4.22)
$\|\boldsymbol{\rho}^{(+)}\|_{1}$ $\leq\zeta^{(+)}_{\max}\,\|\mathbf{V}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{2}^{2}\,\|\mathbf{V}^{\dagger}_{M_{2}}(\boldsymbol{z}^{(+)})\|_{\mathrm{F}}^{2},$ (4.23)
where $\zeta_{\max}^{(-)}:=\max_{j=1,\ldots,M_{1}}\zeta_{j}^{(-)}$ and $\zeta_{\max}^{(+)}:=\max_{j=1,\ldots,M_{2}}\zeta_{j}^{(+)}$.
Let us now consider the case of structured perturbations. Define further matrices
$\mathbf{\mathcal{S}}^{(-)}_{M_{1},2M_{1}}:=\big{(}S_{n,m}^{(-)}\big{)}_{n,m=1}^{M_{1},2M_{1}}\in\mathbb{C}^{M_{1}\times 2M_{1}},\quad\mathbf{\mathcal{S}}^{(+)}_{M_{2},2M_{2}}:=\big{(}S_{n,m}^{(+)}\big{)}_{n,m=1}^{M_{2},2M_{2}}\in\mathbb{C}^{M_{2}\times 2M_{2}},$
and vectors
$(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{1}})^{(-)}$ $:=\left((S^{(n)}_{1})^{(-)},\ldots,(S^{(n)}_{2M_{1}})^{(-)}\right)^{T}\in\mathbb{C}^{2M_{1}}\quad\text{for}\quad n=1,\ldots,M_{1},$
$(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{2}})^{(+)}$ $:=\left((S^{(n)}_{1})^{(+)},\ldots,(S^{(n)}_{2M_{2}})^{(+)}\right)^{T}\in\mathbb{C}^{2M_{2}}\quad\text{for}\quad n=1,\ldots,M_{2},$
by their components
$S_{n,m}^{(-)}=\left(S^{(n)}_{m}\right)^{(-)}:=$ $\frac{1}{\gamma_{j_{n}}^{(-)}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{1}\}}^{\min\{M_{1},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-z_{j_{n}}^{(-)}\,\sum\limits_{k=\max\{1,m-M_{1}+1\}}^{\min\{M_{1},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{1}}(\boldsymbol{z}^{(-)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}$ (4.24)
and
$S_{n,m}^{(+)}=\left(S^{(n)}_{m}\right)^{(+)}:=$ $-\frac{1}{\gamma_{j_{n}}^{(+)}\,z_{j_{n}}^{-2}}\bigg{(}\,\sum\limits_{k=\max\{1,m-M_{2}\}}^{\min\{M_{2},m-1\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k,n}$
$-\frac{1}{z_{j_{n}}^{(+)}}\,\sum\limits_{k=\max\{1,m-M_{2}+1\}}^{\min\{M_{2},m\}}\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{k,n}\,\bigg{(}\left(\mathbf{V}^{T}_{M_{2}}(\boldsymbol{z}^{(+)})\right)^{\dagger}\bigg{)}_{m-k+1,n}\bigg{)}.$ (4.25)
Formulas (4.24) and (4.25) are obtained using similar ideas from Subsection 3.2. In (4.25), we take into account (4.18).
Now we define a structured sensitivity of the pole $z^{(-)}_{j_{n}}$ located inside the unit circle with respect to the Fourier coefficient $\hat{r}_{-m}$ by $\eta_{n,m}^{(-)}:=|S_{n,m}^{(-)}|$ for $n=1,\ldots,M_{1}$ and $m=1,\ldots,2M_{1}$. Similarly, by $\eta_{n,m}^{(+)}:=|S_{n,m}^{(+)}|$ we define a structured sensitivity of the pole $z^{(+)}_{j_{n}}$ located outside the unit circle with respect to the Fourier coefficient $\hat{r}_{m}$ for $n=1,\ldots,M_{2}$ and $m=1,\ldots,2M_{2}$. Note that Fourier coefficients $\hat{r}_{-m}$ and $\hat{r}_{m}$ are $m$th measurements in this case. We define also two vectors
$\boldsymbol{\eta}^{(-)}:=(\eta_{1}^{(-)},\ldots,\eta_{M_{1}}^{(-)})^{T}\in\mathbb{R}^{M_{1}},\quad\boldsymbol{\eta}^{(+)}:=(\eta_{1}^{(+)},\ldots,\eta_{M_{2}}^{(+)})^{T}\in\mathbb{R}^{M_{2}},$
where the entries
$\eta_{n}^{(-)}:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{1}})^{(-)}\|_{2}=\sqrt{(\eta^{(-)}_{n,1})^{2}+\ldots+(\eta^{(-)}_{n,2M_{1}})^{2}},\quad n=1,\ldots,M_{1},$ (4.26)
and
$\eta_{n}^{(+)}:=\|(\boldsymbol{\mathcal{S}}^{(n)}_{2M_{2}})^{(+)}\|_{2}=\sqrt{(\eta^{(+)}_{n,1})^{2}+\ldots+(\eta^{(+)}_{n,2M_{2}})^{2}},\quad n=1,\ldots,M_{2},$ (4.27)
are structured sensitivities for the poles $z_{j_{n}}^{(-)}$ and $z_{j_{n}}^{(+)}$ respectively.
5 Numerical experiments
In this section, we present a few numerical experiments. The algorithm is implemented in Matlab and use IEEE standard floating point arithmetic with double precision. We use the notations $\tilde{z}_{j}$ and $\tilde{\gamma}_{j}$ for the poles and the coefficients, respectively, reconstructed by the method considered in Subsection 4.1. We define the errors for the poles $z_{j}$ and the coefficients $\gamma_{j}$ by the following formulas
[…]
Note that in the case when we deal with noisy data, we assume that a number of poles $M$, as well as numbers of poles located inside and outside the unit circle $M_{1}$ and $M_{2}$ respectively are known.
Table 1: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.1
$\sigma=10^{-3}$ $N=2$ $N=10$
$z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
min $2.54$e–$05$ $2.06$e–$04$ $3.49$e–$06$ $5.56$e–$05$
max $2.85$e–$04$ $3.34$e–$03$ $2.7$e–$04$ $3.15$e–$03$
average $1.34$e–$04$ $2.03$e–$03$ $1.29$e–$04$ $1.06$e–$03$
$\sigma=10^{-1}$ $N=10$ $N=30$
$z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$ $z_{1}^{(-)}=-0.1$ $z_{1}^{(+)}=-2.1$
min $3.2$e–$03$ $4.09$e–$02$ $7.19$e–$03$ $4.15$e–$02$
max $2.28$e–$02$ $3.65$e–$01$ $2.72$e–$02$ $1.7$e–$01$
average $1.09$e–$02$ $1.57$e–$01$ $1.36$e–$02$ $1.05$e–$01$
We employ $2N$ (noisy) Fourier coefficients $\hat{r}_{-k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(-)}$, and $2N$ noisy Fourier coefficients $\hat{r}_{k}(1+n_{k})$, $k=1,\ldots,2N$, to reconstruct poles $z_{j}^{(+)}$. By $(n_{k})_{k=1}^{2N-1}$ we denote the random noise that is generated in Matlab by $\sigma$*randn(2N), i.e., we take multiplicative noise drawn from the standard normal distribution. We take $L=N$ and consider different values of $N$ and $\sigma$.
To compute sensitivities $\rho_{j}^{(-)}$, $\rho_{j}^{(+)}$ and ,$\eta_{j}^{(-)}$, $\eta_{j}^{(+)}$, we use formulas (4.13), (4.19) and (4.26), (4.27) respectively. Upper bounds for sensitivities $\rho_{j}^{(-)}$ and $\rho_{j}^{(+)}$ are computed via (4.14) and (4.20) respectively.
[…]
$\sigma=10^{-3}$ $N=2$ $N=10$ $\sigma=10^{-1}$ $N=10$ $N=30$
min $1.78$e–$04$ $1.17$e–$04$ min $4.03$e–$02$ $2.77$e–$02$
max $2.11$e–$03$ $2.46$e–$03$ max $3.04$e–$01$ $1.44$e–$01$
average $1.44$e–$03$ $8.48$e–$04$ average $1.11$e–$01$ $9.55$e–$02$
Example 5.1.
Let a rational function $r(z)$ in (1.1) be given by parameters $M=2$ and
$\boldsymbol{z}=(-0.1,-2.1),\quad\boldsymbol{\gamma}=(0.5,0.5)$
(see [5, 21]). When $\sigma=0$, i.e. in case of recovery from exact data, we get the following reconstruction errors with $N=2$,
[…]
For $\sigma=10^{-3}$ and $\sigma=10^{-1}$, we compute 10 iterations and present the results in Table 1. Since we are interested to investigate which pole is more sensitive to perturbations, we compute reconstruction errors for poles $z_{1}^{(-)}=-0.1$ and $z_{1}^{(+)}=-2.1$ separately. Results for recovery of the coefficient vector are presented in Table 2.
Table 3: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.2 for $\sigma=10^{-3}$
$N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
min $7.08$e–$05$ $4.31$e–$06$ $2.79$e–$06$ $9.06$e–$06$
max $6.74$e–$04$ $3.17$e–$04$ $1.65$e–$04$ $1.45$e–$04$
average $3.18$e–$04$ $1.48$e–$04$ $7.87$e–$05$ $5.57$e–$05$
$N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
min $1.71$e–$05$ $1.13$e–$06$ $1.31$e–$05$ $8.74$e–$06$
max $2.86$e–$04$ $2.7$e–$04$ $1.89$e–$04$ $1.9$e–$04$
average $1.44$e–$04$ $1.19$e–$04$ $1.04$e–$04$ $9.43$e–$05$
Analyzing numerical results in Table 1, we conclude that for $\sigma=0.1$ our method still reconstructs location of poles quite good. On the other hand, using more (noisy) Fourier coefficients also gives us better recovery results. We notice that a pole $z_{1}^{(-)}=-0.1$ is less sensitive to perturbations than a pole $z_{1}^{(+)}=-2.1$. This numerical observation is also confirmed with sensitivities values. For a pole $z_{1}^{(-)}=-0.1$, we have $\rho_{1}^{(-)}=2\cdot 10^{-1}$ and $\eta_{1}^{(-)}=1.97$, and for a pole $z_{1}^{(+)}=-2.1$, we have $\rho_{1}^{(+)}=9.52\cdot 10^{-1}$ and $\eta_{1}^{(+)}=6.65$. Both unstructured and structured sensitivities are larger for a pole $z_{1}^{(+)}=-2.1$ located outside the unit circle. We compute also vectors of structured perturbations
$(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(-)}=(2.0\cdot 10^{-1},2.0)^{T}\text{ and }(\boldsymbol{\mathcal{S}}^{(1)}_{2})^{(+)}=(4.2,8.82)^{T}.$
That means that a pole $z_{1}^{(-)}=-0.1$ is the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-2}$, and a pole $z_{1}^{(+)}=-2.1$ with respect to perturbation of the Fourier coefficient $\hat{r}_{2}$.
Table 4: Reconstruction errors $|z-z_{j}|$ for each poles $z_{j}$ for Example 5.2 for $\sigma=10^{-1}$
$N=8$ $z_{1}^{(-)}=0.9$ $z_{2}^{(-)}=-0.9$ $z_{3}^{(-)}=0.9\mathrm{i}$ $z_{4}^{(-)}=-0.9\mathrm{i}$
min $6.41$e–$03$ $2.65$e–$03$ $5.14$e–$04$ $1.38$e–$03$
max $2.02$e–$01$ $2.37$e–$02$ $2.89$e–$02$ $7.51$e–$03$
average $3.99$e–$02$ $1.13$e–$02$ $1.06$e–$02$ $1.82$e–$02$
$N=8$ $z_{1}^{(+)}=1.1$ $z_{2}^{(+)}=-1.1$ $z_{3}^{(+)}=1.1\mathrm{i}$ $z_{4}^{(+)}=-1.1\mathrm{i}$
min $8.17$e–$04$ $8.71$e–$04$ $1.93$e–$03$ $2.83$e–$03$
max $2.95$e–$02$ $2.61$e–$02$ $2.48$e–$02$ $2.41$e–$02$
average $1.56$e–$02$ $1.4$e–$02$ $1.24$e–$02$ $1.15$e–$02$
Example 5.2.
We consider an example when poles are located very close to the unit circle. Let a rational function $r(z)$ of type $(7,8)$ in (1.1), i.e with $M=8$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.9,-0.9,0.9\mathrm{i},-0.9\mathrm{i},1.1,-1.1,1.1\mathrm{i},-1.1\mathrm{i}),\quad\boldsymbol{\gamma}=(1,2,3,4,5,6,7,8),$
respectively. We consider first recovery from the exact data, i.e for $\sigma=0$. For $N=8$, we get the following reconstruction errors
$e(\boldsymbol{z})=1.45\cdot 10^{-15},\quad e(\boldsymbol{\gamma})=5.48\cdot 10^{-14}.$
[…]
$N=8$ $\sigma=10^{-3}$ $\sigma=10^{-1}$
min $4.2$e–$03$ $4.25$e–$01$
max $1.73$e–$02$ $1.22$e–$00$
average $9.7$e–$03$ $6.35$e–$01$
In case of recovery from noisy data, we compute 10 iterations and present recovery result for each pole $z_{j}$ in Tables 3 and 4. Analyzing numerical results, we conclude that a pole $z_{1}^{(-)}=0.9$ is the most sensitive to perturbations. This conclusion is confirmed with sensitivities presented in Table 6. Results for recovery of the coefficient vector are presented in Table 5.
Table 6: Sensitivities for Example 5.2.
$j$     poles $z_{j}^{(-)}$      $\eta_{j}^{(-)}$      $\rho_{j}^{(-)}$      bound $\rho_{j}^{(-)}$
1 $0.9$ 0.243 7.748 13.548
2 $-0.9$ 0.121 3.874 6.774
3 $0.9\mathrm{i}$ 0.081 2.582 4.516
4 $-0.9\mathrm{i}$ 0.06 1.937 3.387
$j$     poles $z_{j}^{(+)}$      $\eta_{j}^{(+)}$      $\rho_{j}^{(+)}$      bound $\rho_{j}^{(+)}$
1 $1.1$ 0.056 3.533 5.695
2 $-1.1$ 0.047 2.944 4.745
3 $1.1\mathrm{i}$ 0.04 2.523 4.067
4 $-1.1\mathrm{i}$ 0.035 2.208 3.559
Example 5.3.
Finally, let a rational function $r(z)$ of type $(3,4)$ in (1.1), i.e with $M=4$, be given by vectors of poles and coefficients
$\boldsymbol{z}=(0.2,0.5,2,50),\quad\boldsymbol{\gamma}=(1,1,1,1),$
respectively. For $N=4$ and $\sigma=0$, we get the following reconstruction errors
$e(\boldsymbol{z})=1.42\cdot 10^{-13},\quad e(\boldsymbol{\gamma})=9.27\cdot 10^{-15}.$
Table 7: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-7}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $1.58$e–$08$ $2.9$e–$08$ $6.07$e–$09$ $1.75$e–$03$
max $5.1$e–$07$ $2.49$e–$07$ $2.7$e–$07$ $3.21$e–$01$
average $2.48$e–$07$ $1.3$e–$07$ $1.45$e–$07$ $1.34$e–$01$
Table 8: Reconstruction errors $|z-z_{j}|$ for each pole $z_{j}$ for Example 5.3 for $\sigma=10^{-5}$
$N=4$ $z_{1}^{(-)}=0.2$ $z_{2}^{(-)}=0.5$ $z_{1}^{(+)}=2$ $z_{2}^{(+)}=50$
min $4.24$e–$06$ $1.18$e–$06$ $2.57$e–$06$ $3.16$e–$00$
max $6.45$e–$05$ $2.77$e–$05$ $2.92$e–$05$ $9.46$e+$01$
average $2.31$e–$05$ $1.13$e–$05$ $1.5$e–$05$ $2.02$e+$01$
Then we consider recovery from noisy data with $\sigma=10^{-7}$ and $\sigma=10^{-5}$. We compute 10 iterations and present the corresponding results in Tables 7 and 8. Analyzing numerical results, we can conclude that the pole $z_{2}^{(+)}=50$ is the most sensitive to perturbations. For $\sigma=10^{-5}$ the method does not reconstruct location of this pole. This numerical observation is also predicted by the sensitivities values presented in Table 9. Both unstructured $\rho_{2}^{(+)}=2.204\cdot 10^{3}$ and structured $\eta_{2}^{(+)}=1.577\cdot 10^{4}$ sensitivities are sufficiently larger than sensitivities of other poles. Note, that locations of poles $z_{1}^{(-)}=0.2$, $z_{2}^{(-)}=0.5$ and $z_{1}^{(+)}=2$ are reconstructed even for $\sigma=10^{-1}$. We present results regarding recovery of the coefficient vector in Table 10.
Table 9: Sensitivities for Example 5.3.
$j$     poles $z_{j}^{(-)}$      $\eta_{j}^{(-)}$      $\rho_{j}^{(-)}$      bound $\rho_{j}^{(-)}$
1 $0.2$ $1.807\cdot 10^{1}$ $1.766\cdot 10^{1}$ $3.937\cdot 10^{1}$
2 $0.5$ $1.518\cdot 10^{1}$ $2.249\cdot 10^{1}$ $5.625\cdot 10^{1}$
$j$     poles $z_{j}^{(+)}$      $\eta_{j}^{(+)}$      $\rho_{j}^{(+)}$      bound $\rho_{j}^{(+)}$
1 $2$ $1.973\cdot 10^{1}$ 5.43 $1.993\cdot 10^{1}$
2 $50$ $1.577\cdot 10^{4}$ $2.204\cdot 10^{3}$ $6.477\cdot 10^{3}$
We compute also vectors of structured perturbations for poles inside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(-)}=(0.555,4.999,13.333,11.111)^{T},\quad(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(-)}=(0.222,2.666,9.999,11.111)^{T},$
and outside the unit circle
$(\boldsymbol{\mathcal{S}}^{(1)}_{4})^{(+)}=(3.472\cdot 10^{-3},0.354,9.375,17.36)^{T},$
and
$(\boldsymbol{\mathcal{S}}^{(2)}_{4})^{(+)}=(54.25,2.929\cdot 10^{3},1.106\cdot 10^{4},1.085\cdot 10^{4})^{T}.$
[…]
$N=4$ $\sigma=10^{-7}$ $\sigma=10^{-5}$
min $7.38$e–$04$ $3.79$e–$02$
max $7.38$e–$03$ $6.95$e–$01$
average $4$e–$03$ $4.11$e–$01$
That means that poles $z_{1}^{(-)}=0.2$ and $z_{2}^{(-)}=0.5$ are the most sensitive with respect to perturbations of the Fourier coefficient $\hat{r}_{-3}$, the pole $z_{1}^{(+)}=2$ – the Fourier coefficient $\hat{r}_{4}$, and the pole $z_{2}^{(+)}=50$ – the Fourier coefficient $\hat{r}_{3}$.
Table 11: $\ell_{2}$-norms and upper bounds of $\ell_{2}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$\|\boldsymbol{\rho}^{(-)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{2}$) $\|\boldsymbol{\rho}^{(+)}\|_{2}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{2}$)
$N=4$ $2.86$e+$01$ $6.866$e+$01$ $2.204$e+$03$ $6.477$e+$03$
Table 12: $\ell_{1}$-norms and upper bounds of $\ell_{1}$-norms of vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ for Example 5.3
$\|\boldsymbol{\rho}^{(-)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(-)}\|_{1}$) $\|\boldsymbol{\rho}^{(+)}\|_{1}$ bound ($\|\boldsymbol{\rho}^{(+)}\|_{1}$)
$N=4$ $4.01$e+$01$ $5.725$e+$01$ $2.209$e+$03$ $6.802$e+$03$
Further we compute the $\ell_{1}$- and $\ell_{2}$-norms of the vectors $\boldsymbol{\rho}^{(-)}$ and $\boldsymbol{\rho}^{(+)}$ and their upper bounds obtained in (4.21), (4.22) and (4.23). We present the results in Tables 11 and 12. Analyzing these results, we conclude that for this example upper bounds (4.21), (4.22) and (4.23) are quite sharp.
Acknowledgement
The author gratefully acknowledge support by the German Research Foundation in the framework of the RTG 2088. The author would like to thank Ion Victor Gosea (MPI Magdeburg) for several valuable discussions regarding sensitivities of the eigenvalues of a matrix pencil problem.